Exercise 16-18

Swifty Inc. presented the following data.
Net income $2,340,000
Preferred stock: 53,000 shares outstanding, $100 par, 8% cumulative, not convertible 5,300,000
Common stock: Shares outstanding 1/1 692,400
   Issued for cash, 5/1 321,600
   Acquired treasury stock for cash, 8/1 160,800
   2-for-1 stock split, 10/1

Compute earnings per share. (Round answer to 2 decimal places, e.g. $2.55.)
Earnings per share $
Answer #1
Statement of Total Shares Out standing
Converted to shares Shares 2 for 1 stock Total Shares Total Shares
Additional shares
comman stock (1/1) 692400 (692400/1=692400) (692400/2)=346200 (692400+346200) 1038600
Common stock (5/1) 321600 (321600/5=64320) (64320/2)=32160 (64320+32160) 96480
Common stock (8/1) 160800 (160800/8=20100) (201000/2)=10050 (21100+10050) 30150
Total Shares Outstanding 1165230
Net Income 2340000
Preference Dividend (5300000 * 8%) 424000
Net income Avilable for Share Holders 1916000
EPS = Net Income available for Share holders 1916000 = 1.64
Equity share outstanding 1165230